OpenStudy (anonymous):

How would you expand 4(x+7)

OpenStudy (mathteacher1729):

You would distribute: 4*x + 4*7. "Four times x plus four times 7".

OpenStudy (anonymous):

You the distrubutive property

OpenStudy (anonymous):

you need to multiply the outside number with the numbers inside the paranthese

OpenStudy (anonymous):

I don't know, sorry I'm very very very very bad at math

OpenStudy (solomonzelman):

For example \(\large\color{black}{ \displaystyle \color{red}{2}(\color{green}{h}+\color{blue}{6}) }\) \(\large\color{black}{ \displaystyle (\color{red}{2}\times \color{green}{h})+(\color{red}{2}\times \color{blue}{6}) }\) \(\large\color{black}{ \displaystyle \color{red}{2}\color{green}{h}+\color{black}{12}) }\)

OpenStudy (solomonzelman):

(the very last parenthesis at the end is a typo)

OpenStudy (anonymous):

sooooooo the answer would be??

OpenStudy (mathteacher1729):

The answer is: 4*x + 4*7. "Four times x plus four times seven". Only instead of writing "4*7" you actually write whatever that product is. :) That's all. Nothing more do here.